Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20120917

Observers : Iair Arcavi, Yen-Chen Pan. 
Data reduction team : K. Maguire, M. Sullivan, A. Gal-Yam, O. Yaron

Clouds and high humidity.
*** Closed from sunset until 00:30 UT (one hour after -12 degrees) ***
EFOSC spec standards: EG274, Feige110
SN2009ip [EFOSC: gm11 gm16] - Followup
Seeing ~2". Clouds. Aborted after ~15 minutes into gm11 due to rising humidity, but got something.
*** Closed from 1:15 UT until 6:45 UT ***
[SN2012ec skipped because done at Asiago.]
R71 [EFOSC: UBVRi gm11 gm16] - Followup
Seeing ~2".
SN2012Z [EFOSC: gm13] - Followup
Seeing ~2.2". Aborted after ~30 minutes due to clouds, but got something. Ca-rich?
*** Closed from 8:15 UT until sunrise ***
Comments